For which values of a and b is the line b 2x + y = tangent to the parabola 2

y = ax

when x = 2? (2) 

 


1
Expert's answer
2021-02-24T07:02:39-0500

"\displaystyle
y = ax^2 \\

y' = 2ax \\

y'_{\{\textsf{at}\,\, x = 2\}} = \textsf{Gradient} = 2(2)(a) = 4a \\

\textsf{At}\,\,\, x = 2, y = 4a \\

\textsf{Equation of tangent}\\

y - 4a = 4a(x - 2)\\

y = 4ax + 4a - 8a = 4a(x - 1)\\

\textsf{If the line is equal with}\,\,2x + y = b,\,\,\textsf{we can compare}\\

2x + y = b \\

y - 4ax = -4a\\

2 = -4a, a = -\frac{1}{2},\,\, b = -4a = -4 \times -\frac{1}{2} = 2\\

\therefore a = -\frac{1}{2},\,\,\textsf{and}\,\, b = 2"